OpenStudy (anonymous):

SOLVE FOR Y: x = 7y x + y + z = 80 2x - y + z = 105 A. -7 B. 15 C. -175/13 D. 35 E. none of these

OpenStudy (danjs):

ok

OpenStudy (anonymous):

This one i think is D? i know you have to solve by substitution.

OpenStudy (danjs):

Take the first equation x = 7y, and replace x in the second equation with that

OpenStudy (danjs):

7y + y + z = 80 2(7y) - y + z = 105 I replaced X in the second 2 equations with x= 7y, good with that so far?

OpenStudy (anonymous):

yes.

OpenStudy (danjs):

now you have 2 equations with 2 unknowns Y and Z,, we can solve this by a couple methods

OpenStudy (danjs):

adding the Y's to simplify a little bit 8Y + Z = 80 13Y + Z = 105

OpenStudy (danjs):

we can multiply one of the equations by a constant number and then add them together,

OpenStudy (danjs):

here, if you multiply one of them by -1, then when you add them you get +Z-Z = 0 the Z will go to zero

OpenStudy (danjs):

(-1)(8Y + Z = 80 ) multiply first by -1 13Y + Z = 105 ------------------- -8Y - Z = -80 13Y + Z = 105 ADD the two together -------------- 5Y + 0Z = 25 no Z's Left now 5Y = 25 divide by 5 both sides Y = 5

OpenStudy (danjs):

Do you recall doing this kind of thing in class?

OpenStudy (anonymous):

okay that's what i was getting as well, i got 35 for x 5 for y and 40 for z when i did it.

OpenStudy (anonymous):

yes we did this in class.

OpenStudy (danjs):

ok, just want to make sure i am using a similar method to what you are learning...

OpenStudy (danjs):

Now that you figured Y=5, you can use that in the first equation to get X X = 7Y = 7*5 = 35 X=35

OpenStudy (danjs):

Now you know X and Y, use any of the last 2 equations to get Z

OpenStudy (anonymous):

which would come to z = 40, right?

OpenStudy (danjs):

Using equation 2, x + y + z = 80 and X=5 , Y = 35 5 + 35 + z = 80 z = 40 yep

OpenStudy (anonymous):

alright :) i'm glad i'm doing this correctly then. so the right answer would be E for this then. you're the best!! would you mind helping me with the other one as well?

OpenStudy (danjs):

If you wanted to check after you finished, use x=5 y=35 z=40 plug those into all 3 equations and see if they all are true

OpenStudy (danjs):

oh, Y=5, X=35, Z=40. typo above

OpenStudy (anonymous):

okay! thank you. im glad that i know this method correctly.

OpenStudy (danjs):

Right , answer is NONE, Y=5

OpenStudy (danjs):

cool, welcome, anytime